# On the Surprising Effectiveness of Spectral Clipping in Learning Stable Linear Dynamics

### v4

@@ 6 Conclusion
Appendix A Proof for Arbitrarily Small System Evolution Error
@@
For any $\eta>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, we can always find its diagonalizable counterpart $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
Proof.
@@
where $E\in\mathbb{R}^{n\times n}$ is the residual matrix and, from Lemma. 3.1, $||E||_{1}$ can be arbitrarily small (i.e., $||E||_{1}<\gamma$). Then we have
@@
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}=||\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}<=\sum_{i=1}^{k}||{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}$ (6)
@@
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<m\gamma$ (7)
Let $\eta=m\gamma$. Note that since Lemma. 3.1 guarantees the existence of a diagonalizable matrix $\tilde{A}_{LS}$ that satisfies an arbitrarily small $\gamma$, it follows that $\eta=m\gamma$ can also be made arbitrarily small. Therefore, for any $\eta>0$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$. ∎
Appendix B Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct two different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see [13]), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in [47].
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m [13], and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$ (this criterion is the same as that used in [13]), and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
Appendix C Spectrum Clipping for Controlled System
@@
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n},\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ are the system state and control signal at time $t$, respectively. Similarly, $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, and $B\in\mathcal{S}_{B}$ is the unknown control matrix, and $\mathcal{S}_{B}\subset\mathbb{R}^{n\times m}$.
@@
Appendix D Qualitative performance on the Franka Panda dataset
From the results shown in Fig. 9, we can observe that CG and SOC overall exhibit the best performance, especially when using only 100 training samples. However, as dataset size increases, the performance of SC is significantly improved, and become comparable to both CG and SOC, while still being orders of magnitude faster for training.
Figure 9: Rollouts of the end effector trajectory. Each subplot contains 25 trajectories, representing 5 different initial positions with matrices trained on 5 subsets.
Appendix E Tradeoff between stability and expressivity
As mentioned in Section. 3, we can also select the $\varepsilon$ values to ensure that the resulting linear system meets the criteria for well-known Lyapunov stability. Through the experiments conducted on UCSD, DTDB, DexManip, and Franka Panda datasets, we demonstrate that this strategy does not only guarantee the Lyapunov stability, but also achieve comparable performance, provided that an appropriate $\varepsilon$ is chosen, such as $\varepsilon=10^{-5}$. Note that we omit the computation time and memory usage results, as this strategy does not impact these metrics.
For UCSD and DTDB dataset, we first compute the Reconstruction error of matrices clipped to different $\varepsilon$ values,
@@ Appendix E Tradeoff between stability and expressivity
For DexManip dataset, the performance of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are also nearly distinguishable, as shown in Fig. 15 and Fig. 16 (Reconstruction error) and Table. 8 and Table. 9 (Success & Safety rates). Also, we can observe that $SC(\varepsilon=10^{-2})$ lacks the expressivity needed to generate diverse motions, leading to a significant low task success rate, although it maintains a 100% safety rate as all states converge to zero.
For Franka Panda dataset, though SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ still perform similarly, we observe that $SC(\varepsilon=10^{-2})$ also performs similarly, and even performs the best when using the fewest data points for training (see Fig. 17). We conjecture that this is due to the control inputs and feedback controller. In addition, all these three performs comparably in the long-term prediction (Table. 10).
Figure 12: DTDB Sequence 185: A flushing toilet.
Figure 13: DTDB Sequence 205: A running stream.
Figure 14: Qualitative results of 2 sequences from DTDB dataset (same sequences as used in Fig. 6). Matrices are clipped to 1, clipped to $1-\varepsilon$, or not clipped.
@@
SC($\varepsilon=10^{-2}$) 706.0687 inf 0%
[…]
@@
Tool Tool_short Reloc Reloc_with_failures
LS 55.5% 3.3% 96.0% 9.6%
SC($\varepsilon=0$) 94.0% 90.7% 93.8% 94.4%
SC($\varepsilon=10^{-5}$) 94.0% 89.0% 94.4% 93.8%
SC($\varepsilon=10^{-2}$) 24.7% 0% 66.7% 42.9%
@@
Tool Tool_short Reloc Reloc_with_failures
LS 0% 0% 100.0% 0.0%
SC($\varepsilon=0$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-5}$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-2}$) 100.0% 100.0% 100.0% 100.0%
@@
Figure 15: Average reconstruction error on the Tool and Reloc datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Figure 16: Average reconstruction error on the Tool_short and Reloc_with_failures datasets datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.

### v6

@@ 6 Conclusions
Appendix A Approximation error bounds for non-diagonalizable system matrices
Lemma A.1.
For any $\gamma>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. $\hat{A}_{LS}-\tilde{A}_{LS}||_{1}<\gamma$.
Proof A.2.
Lemma. A.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory (Horn and Johnson, 2012; Trefethen and Bau, 2022).
@@ Appendix A Approximation error bounds for non-diagonalizable system matrices
For any $\eta>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, we can always find its diagonalizable counterpart $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
Proof A.4.
@@
where $E\in\mathbb{R}^{n\times n}$ is the residual matrix and, from Lemma. A.1, $E||_{1}$ can be arbitrarily small (i.e., $E||_{1}<\gamma$). Then we have
@@
[…]
@@
$\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<m\gamma$ (7)
Let $\eta=m\gamma$. Note that since Lemma. A.1 guarantees the existence of a diagonalizable matrix $\tilde{A}_{LS}$ that satisfies an arbitrarily small $\gamma$, it follows that $\eta=m\gamma$ can also be made arbitrarily small. Therefore, for any $\eta>0$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
Appendix B Spectrum Clipping for Controlled Systems
@@
[…]
@@
The experiment results can be found in Appendix. E.
Appendix C Tradeoff between stability and expressivity
As mentioned in Section. 3, the choice of $\varepsilon$ values establishes a trade-off between expressivity and stability. In the experiments reported below, we systematically analyze this fundamental trade-off. We omit the analysis of compute and memory efficiency since $\varepsilon$ does not impact them.
UCSD and DTDB dataset: We first compute the Reconstruction error of matrices clipped to different $\varepsilon$ values,
@@ Appendix C Tradeoff between stability and expressivity
Figure 7: Average reconstruction error as a function of subspace dimensions of the UCSD dataset.
@@
DTDB Sequence 185: A flushing toilet.
DTDB Sequence 205: A running stream.
Figure 8: Qualitative prediction results on two example sequences from DTDB dataset.
@@
SC($\varepsilon=10^{-2}$) 706.0687 10000.0 0%
Training Set 18.8768 (average over all training sequences) 100.0%
@@
DexManip dataset: We find that the performance of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are also nearly distinguishable, as shown in Fig. 9 (Reconstruction error) and Table. 6 and Table. 7 (Success & Safety rates). Also, we can observe that $SC(\varepsilon=10^{-2})$ lacks the expressivity needed to generate diverse motions, leading to a significant low task success rate, although it maintains a 100% safety rate as all states converge to zero.
Figure 9: Average reconstruction error on the Reorient, Reloc, Door, Tool, Reloc_with_failures and Tool_short datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Reorient Door Tool Tool_short Reloc Reloc_with_failures
LS 69.5% 75.3% 55.5% 3.3% 96.0% 9.6%
SC($\varepsilon=0$) 75.6% 76.6% 94.0% 90.7% 93.8% 94.4%
SC($\varepsilon=10^{-5}$) 76.3% 76.6% 94.0% 89.0% 94.4% 93.8%
SC($\varepsilon=10^{-2}$) 83.2% 12.0% 24.7% 0% 66.7% 42.9%
@@
Reorient Door Tool Tool_short Reloc Reloc_with_failures
LS 100.0% 100.0% 0% 0% 100.0% 0.0%
SC($\varepsilon=0$) 100.0% 100.0% 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-5}$) 100.0% 100.0% 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-2}$) 99.2% 100.0% 100.0% 100.0% 100.0% 100.0%
@@
Franka Panda dataset: Though all variants of SC perform similarly with sufficient data, we observe that $SC(\varepsilon=10^{-2})$ is the most sample-efficient and comparable to CG and SOC (see Fig. 10 and Table. 8). This is likely because the true spectral radius of the underlying system is likely closer to $0.99$.
@@
Appendix D Connections between SC and Koopman modes
We have also discovered interesting connections between SC and Koopman modes, as mentioned in Section 5.6, which provides us with valuable insights on how SC works on corrupted datasets. For each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of Koopman matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following (Rowley et al., 2009), we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from Reloc_with_failuresReloc\_with\_failures dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 11, respectively (see the supplementary video for full policy rollouts). It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 4). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 11. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 4). The full policy rollout is also available in the supplementary video.
Appendix E SC can help learn stable LDS with control inputs
We finally evaluated SC’s ability to learn stable LDSs with control inputs (see Appendix. B for ). To this end, we used Franka Panda (Gaz et al., 2019), a dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of (randomly generated) control inputs and the corresponding robot states.
We first evaluate the long-horizon prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs, and then compute the average $\mathscr{l}_{2}$ distance between the predicted and ground-truth end-effector positions for each method. From the results shown in Table. 9, we observe that while SC performs worse than CG and SOC when $d=100$, its performance becomes comparable as $d$ increases. Notably, SC achieves this with significantly lower computation time and memory usage.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in Mamakoukas et al. (2020)). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Fig. 10 of Appendix. C, where SC demonstrates performance comparable to that of other methods.
Dataset size LS CG WLS SOC SC
$d=100$ $0748\pm 0.1218 0.0$$0166\pm 0.0066$ $3390\pm 0.2059 0.0$$225\pm 0.0131 0.0$$0693\pm 0.1370$
$d=2000$ $0207\pm 0.0173 0.0$$0154\pm 0.0058$ $3051\pm 0.1853 0.0$$154\pm 0.0058 0.0$$0183\pm 0.0138$
$d=3000$ $0164\pm 0.0083 0.0$$0155\pm 0.0058$ $2845\pm 0.1501 0.0$$155\pm 0.0058 0.0$$0156\pm 0.0057$
Appendix F Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct four different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see Han et al. (2023)), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.
Pen Reorientation: This task is mainly about orienting a pen to a random target rotation angle. The hand and object states are lifted to a higher-dimensional $z$ states and different approaches are used to learn the LDSs in the $z$ state space. This dataset consists of 131 rollouts that pens are reoriented to the target rotation angles successfully ($T=100$). At each time step $t$, if $\mathrm{o}^{\text{goal}}\cdot\mathrm{o}^{\text{pen}}_{t}>0.90$ ($\mathrm{o}^{\text{goal}}\cdot\mathrm{o}^{\text{pen}}_{t}$ measures orientation similarity), then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T}\rho(t)>10$, and ii) safe if the pen remains in the robot hand at the final time step (i.e., it is not dropped).
Door Opening: This task is mainly about opening a door from a random starting position. The hand and object states are lifted to a higher-dimensional $z$ states and different approaches are used to learn the LDSs in the $z$ state space. This dataset contains 158 rollouts in which the doors are fully opened ($T=70$). The task is considered i) successful if at last time step $T$, the door opening angle is larger than 1.35 rad, and ii) safe if the robot hand is still near the door (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in Rajeswaran et al. (2018a).
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m, and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$, and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.